three forces of 30N, 20N, 10N … are acting alond three sides of equilateral triangle in cyclic order. find the magnitude, direction and position of the resultant.

please solve this question complete.

Grade:12th pass

1 Answers

Shubham Maurya
6 years ago
Forces A,B,C are 30N,20N,10N respectively. Let force A act along+ve x-axis thenForce A = 30iForce B = (20cos120°)i + (20sin120°)jForce C= (10cos240°)i +(10sin240°)jSo,Force A = 30iForce B= -10i +10√3jForce C= -5i -5√3jResultant force(Position)=15i +5√3jMagnitude of resultant = 10√3Direction of resultant = Resultant force vector/magnitude= (√3/2)i + (1/2)jI`ve answered all the questions.Hope you`ve understoodRegardsShubham Maurya,Student of Don Bosco AcademyPatna